In the time it takes one car to travel 93 km, a second car travels 111 km. If the average speed of the second car is 12 km/h faster than the speed of the first car, what is the speed of the car?
Answer by palanisamy(496)

Let the speed of the first car be x km/h
Then the speed of the second car = x+12
Time taken by tne first car to travel 93kms = 93/x
Time taken by tne second car to travel 111 kms = 111/(x+12)
Given, both times are equal.
93/x = 111/(x+12)
93*(x+12) = 111x
93x+1116 = 111x
1116 = 111x-93x
1116 = 18x
x = 1116/18
x = 62
So the speed of the first car is 62 kms/h
and the speed of the second car is 62+12 = 74 kms/h